OpenStudy (anonymous):

Chain rule help? How is y=(1-(1/x))^2 supposed to derive to (2/x^2)-(2/x^3)?

OpenStudy (anonymous):

\[\left(1+\frac{1}{x}\right)^2=1+\frac{1}{x}-\frac{2}{x^2}\] is a start

OpenStudy (anonymous):

if you mean to take the derivative, your answer is not right

OpenStudy (anonymous):

How did you get the subtraction of (2/x^2)?

OpenStudy (anonymous):

multiplied out!

OpenStudy (anonymous):

ok all that and i made a typo \[\left(1-\frac{1}{x}\right)^2=\left(1-\frac{1}{x}\right)\times \left(1-\frac{1}{x}\right)=...\] thats better

OpenStudy (anonymous):

Okay, thanks! Let me try it and see if I get it. (: